EVA Report – March 2nd

Crew 263 EVA Report 02-03-2022

EVA #13

Author: Mathéo Fouchet

Purpose of EVA: Changing the batteries of atmospheric experiments, Testing an EVA security protocol, and Search and Rescue exercise with a Thermal Drone with teams less than a kilometer apart.

Start time: 9:13 am
End time: 10:25 am

Narrative: We left the airlock at 9:13 and checked the rover batteries (all full), then we took Curiosity and Opportunity as requested and went to the atmospheric experiment at 9:16. We changed the batteries of MegaAres and the LOAC and we put back the Purple Air device. We went back to rovers at 9:26 to go to Robert’s Rock Garden, Opportunity and Curiosity had 100% battery.

We noticed that we passed Robert’s Rock Garden for approximately 20 meters, so we turned around and stopped the rovers next to Robert’s Rock Garden at 9:34. We checked the rovers’ batteries: Opportunity at 98% and Curiosity at 95%.

To implement the Search and Rescue with Drone experiment, the Blue team (Valentine and Elena) walked in the East direction. Then, the Red team launched the drone at 9:46 to look for the Blue team. During this time, the Blue team tried to implement the Security Protocol, everything went well and Valentine acted the «victim» well. The Red team saw the Blue team at 9:50 thanks to the drone’s camera. The drone stayed in the air to help the Red team to reach the Blue team, we carefully found an easy way to join the others. Both teams gathered at 9:57 thanks to the drone vision. All 4 EVA members reached the rovers (still 98% for Opportunity and 95% for Curiosity) at 10:17. The radio contact between the two teams has never been lost.

We finally stopped and plugged the rovers at the RAM and reached the airlock at 10:25.

Destination: Intersection between Entrance Road and Cow Dung Road, Zone I, and Kissing Camel Ridge E

Coordinates (use UTM NAD27 CONUS): Atmospheric Experiment: 518288 E, 4250586 N; 4249500W, 518350S (entrance of Robert’s Rock Garden), 4249500W, 518500S (Kissing Camel Ridge E.)

Participants: Valentine Bourgeois, Mathéo Fouchet (EVA Leader), Cerise Cuny, Elena Lopez-Contreras.

Road(s) and routes per MDRS Map: Entrance Road, Cow Dung Road.

Mode of travel: Foot at Kissing Camel Ridge E, Rover until Robert’s Rock Garden

Sol Summary – March 2nd

Crew 263 Sol Summary 02-03-2022

Sol: 9

Summary Title: Mid mission

Author’s name: Cerise Cuny

Mission Status: Nominal

Sol Activity Summary: This morning, the first shower of the sim was taken. Léa only used 1.8 gal of water, which is a very good score compared to normal, non-Martian showers. We will try to all be as economical as her for our own first shower. The EVA was breathtaking. We went to the south to Kissing Camel. I found the landscape of Robert’s Rock Garden absolutely mesmerizing, and it was nice seeing a very different landscape from the surroundings of the Hab.

At lunch, we had debates about the French educational system. This debate has been recurrent during the mission preparation and some of us are getting tired of the same arguments being used again and again. We decided not to talk about it anymore.

In the afternoon, we did some ultrasound image acquisitions for the CNES, and two of us had to go through a very exhausting human factor experiment – one of the tasks is particularly long and tiresome. Marine will try doing some cheese naan for this evening’s dinner, following Clément and Crew 240 recipe. They also gave us a recipe for homemade hamburgers – I can’t wait to try them. Speaking of food, we did a food inventory and compared it to the food inventory we did at the beginning of the mission.

We ate more carbs per person than the serving sizes on the boxes, but I believe we need the energy to do sports and EVAs. We only eat enough food not to be hungry, thus I wouldn’t be comfortable reducing the food servings. We already talked about reducing portions but its provoked anxiety among the crew members.

Anomalies in work: None.

Weather: Clear sky, warm.

Crew Physical Status: Good.

EVA: EVA 13, in the morning, went very well.

Reports to be filed: Sol Summary, EVA Request (n°14), EVA Report (n°13), Operations Report, Journalist Report, HSO Report.

Support Requested: Food Supplies: We won’t have enough cereals and oat for the third week. We also might need a little bit more flour. Wi-Fi: Our Astronomer has an issue with the Astronomy Wi-Fi. The laptop does connect to the Wi-Fi but displays ‘No Internet’.

Operations Report – March 2nd

Crew 263 Operations Report 02-03-2022

SOL: 9

Name of person filing report: Mathéo Fouchet

Non-nominal systems: NA

Notes on non-nominal systems: NA


Spirit rover used: No

Hours: 157.2

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Opportunity rover used: Yes

Hours: 85.1

Beginning charge: 100%

Ending charge: 94%

Currently charging: Yes

Curiosity rover used: Yes

Hours: 169.1

Beginning charge: 100%

Ending charge: 87%

Currently charging: Yes

Perseverance rover used: No

Hours: 228.8

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

General notes and comments: NA

Summary of Hab operations:

WATER USE: 17 gallons

Water (static tank): 361 gallons

Water (loft tank): 38 gallons

Water Meter: 01562097 units

Static to Loft Pump used: Yes

Static tank pipe heater: On

Static tank heater: On

Toilet tank emptied: No and it smells better because we emptied it yesterday

Summary of internet: NA

Summary of suits and radios: Suit #10 has not been used today and was charging all day, but its battery voltage is currently at 9.6V. I was wondering if it would be possible to charge it “manually” or to replace the battery using the spare ones in the deck near the front airlock. Let me know if you have any suggestions.

Summary of GreenHab operations: NA

Summary of ScienceDome operations:

Dual split: Off

SOC: 97% at 6:00pm

Summary of RAM operations: Broom to sweep the RAM

Summary of any observatory issues: NA

Summary of health and safety issues: See HSO Report

Questions, concerns and requests to Mission Support: NA

Journalist Report – March 2nd

Sol 9 – A Hollywoodian movie

Author: Nicolas Wattelle

We are now at a point where the end of our mission is as far as its beginning. It is a strange feeling; this is one of the first times where we concretely think about the “end” of our more than a one-year-long project. But be reassured, it does mean that we do not keep up the good work!

Valentine is a clever strategist. Our HSO, for Health and Security Officer, takes care of us every day with sports sessions, proposing sophrology exercises, keeping an eye on our personal physiological data… She is like the guardian angel of each of us here, and it is great to have her by our side. But this morning, she organised an exercise of rescuing an injured astronaut, but in order to provide great data to analyse, she developed a scenario, Hollywood-worthy.

All began during yesterday’s eve, where she separately took each member of the crew to explain their role in testing a rescue protocol. She told everyone, except Cerise and Léa, that Elena would simulate a fall and an injury, and everyone one would have to abide by the protocol. The aim was to measure different parameters such as efficiency, time to secure the victim, and physiological data from Martianauts living the situation in EVA.

But in reality, she told the true scenario to Léa (who had to stay in the Hab during the EVA) and Cerise (who was Mathéo’s buddy during the EVA): Valentine herself, was meant to simulate being the victim before Elena did. Like this, the situation was more likely unforeseen for most of us and would be more realistic, while keeping the situation under control if something went wrong. In this case, Léa would have been able to inform Marine (today’s Habcom) and same with Cerise and Mathéo. It was very clever.

Everything went as planned, and after the protocol was applied, Valentine revealed the truth to the Marine, Elena, and Matheo who couldn’t be sure of what had really happened (even if Elena seemed to have understood quickly…). And we are now waiting for the data to be processed. This was a very exciting morning, and we had a great time talking about it at lunch!

For the work that has been done inside the station today, we made our second session of ultrasound with the Sonoscanner device and CNES protocol. Since last week, we received some adjustments from researchers, and it was useful. I think everyone improved their “image capture skills”, finding on average one organ more than last time.

On this subject, Elena is really impressive for an engineering student, she made twice a perfect shot! She may begin a conversion… The whole crew seems to appreciate this experiment and is really involved in giving their feedback and feelings about the protocol. This is a pleasure working in these conditions!

Supplemental Operations Report – March 2nd

Supplemental Operations Report 2 Mar 2022

Name of person filing report: Shannon Rupert

Reason for Report: Routine

Non-nominal systems: Nothing to report

Action taken for non-nominal systems: n/a

Generator: Still being manually run at night. Approximate hours 9 pm to 7 am.

ScienceDome Dual Split: Off

Solar—Charge controllers off at night.

Solar— SOC is nominal. Nice to be able to see what’s going on again.

Notes on power system: We are once again back where we were at the beginning of the season. Both the generator and the solar system are nominal. Autostart is not working.

Propane Reading, station tank – 70 %

Propane Reading, director tank— 60 %

Propane Reading, intern tank— 54 %

Propane Reading, generator— 41 %

Ethanol Free Gasoline – not known gallons. There are four gas cans together. Three of them have old gas with ethanol and one doesn’t, but I don’t know which one that is.

Water (Outpost tank) – 100 gallons

Sojourner rover used: No

Hours: 161.6

Beginning charge: 100

Ending charge: 100

Currently charging: Yes
Notes on rovers: Nothing to report

ATV’s Used: (Honda, 350.1, 350.2, 300): Honda, 350 (whichever one is working)

Reason for use: making sure they are in operating order

Oil Added? No.

ATV Fuel Used: 0 gallons

# Hours the ATVs were Used: 0.5 hours

Notes on ATVs: Both started quickly and ran well.

HabCar used and why, where? no

CrewCar used and why, where? no

Luna used and why, where? Yes, to town for package pickup

Campus-wide inspection, if action is taken, what and why? Nothing to report

General notes and comments: Nothing to report

Summary of general operations: Nothing to report

Summary of internet: The Hab account is still online. The technician who is doing the upgrade still does not know when he will be here. I installed Starlink today but need more information about our account before making it operational.

Summary of suits and radios: Nothing to report

Summary of Hab operations: Added a humidifier to the upper deck. We purchased them for the trailers and found the added humidity in the air is very nice.

Summary of Outpost operations: The intern trailer’s greywater tank is broken. I have not looked at it yet.

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Added analog tire gauge to inventory. Currently, it is in the drawer with the water filter wrench, for no reason except that it fit.

Summary of any observatory issues: Waiting for a second part which is scheduled to arrive tomorrow.

Summary of health and safety issues: Nothing to report.

Copyright © The Mars Society. All rights reserved. | Main Site